what is the vertex and axis of symetry of parabola with equation y=2x^2 +8x+3.
I tried to complete the square and i got the vertex to be (-2,-1) but the answer sheet the teacher gave says the vertex is (-2,-5)? What did I do wrong?

Answer by stanbon(75887)

what is the vertex and axis of symetry of parabola with equation y=2x^2 +8x+3.
y-3=2(x^2+4x
Complete the square on the right side and maintain the equality
by adding the same term to he left side, as follows:
y-3+2(4)=2(x^2+4x+4)
y+5=2(x+2)^2
Axis of symmetry is x=-2
Vertex is (-2,-5)
Cheers,
Stan H.